# Humanoid Locomotion and Manipulation: Current Progress and Challenges in Control, Planning, and Learning thanks: *co-corresponding authors

### v1

@@ Abstract
Humanoid robots have great potential to perform various human-level skills. These skills involve locomotion, manipulation, and cognitive capabilities. Driven by advances in machine learning and the strength of existing model-based approaches, these capabilities have progressed rapidly, but often separately. Therefore, a timely overview of current progress and future trends in this fast-evolving field is essential. This survey first summarizes the model-based planning and control that have been the backbone of humanoid robotics for the past three decades. We then explore emerging learning-based methods, with a focus on reinforcement learning and imitation learning that enhance the versatility of loco-manipulation skills. We examine the potential of integrating foundation models with humanoid embodiments, assessing the prospects for developing generalist humanoid agents. In addition, this survey covers emerging research for whole-body tactile sensing that unlocks new humanoid skills that involve physical interactions. The survey concludes with a discussion of the challenges and future trends.
@@ I Introduction
Figure 1: Humanoids executing locomotion and manipulation tasks: (a) HRP-4 wipes a wood board while adapting to terrain [1]; (b-g) Object pick and place by Digit, Hector [2], Atlas, H1, Justin [3], and Apollo; (h) iCub pushes a cart [4]; (i) Nadia opens a door [5]; (j-k) Object manipulation by Figure 01 and Optimus; (l) MIT humanoid whole-body push recovery [6].
Humanoid robots are well suited for executing human-level tasks, as they are built to (ideally) replicate human motions in achieving various whole-body loco-manipulation tasks, e.g., applications ranging from manufacturing to services, as shown in Fig. 1. Their anthropomorphism makes them stand out from other robot forms in terms of these human-like tasks. Humanoid robots can interact with humans for physical collaboration tasks, such as collaboratively moving a heavy and large table upstairs and human assistance. However, simultaneously achieving these intricate tasks while addressing highly complex robot dynamics is still challenging, let alone safe physical collaborations with humans and/or operations in unstructured environments. As a promising direction to solve this problem, humanoids could exploit the abundance of data available for and/or from humans to quickly acquire motor and cognitive skills. Therefore, leveraging human knowledge for humanoid embodiment is potentially a fast route to embodied intelligence.
Cognitive and autonomous capabilities in robotics at large are thriving at an unprecedented pace. Perception algorithms can detect, classify, and segment a wide variety of objects in real time. Model-based methods that leverage predictive control and reactive control have enabled agile and reliable locomotion and manipulation. Meanwhile, deep learning policies have demonstrated convincing control results on robot hardware through exploration and imitation. Large foundation models trained on massive, internet-scale datasets began to show capabilities of open-world reasoning. Consequently, the building of autonomous humanoid robots for real-world applications has become possible, leading to the emergence of many humanoid robot companies and concrete deployment applications. Especially with powerful GPU-based parallelization capabilities, companies such as NVIDIA and companies with physical humanoid technologies, such as Boston Dynamics, Tesla, and Figure began collaboration on the embodied intelligence of humanoid robots.
@@
Acknowledging the rapid advancements in humanoid robotics, this article reviews recent developments in Humanoid Locomotion and Manipulation (HLM). As laid out in Fig. 2, humanoid robotics is a multidisciplinary field that spans domains in design, actuation, sensing, control, planning, and decision-making. In this survey, we mainly examine task planning, motion planning, policy learning, and control from the perspective of model-based methods and learning-based methods. Each of these topics has extensive studies, and we aim to spotlight representative works within each topic. For each section, we provide survey papers for further reading. We first synthesize traditional model-based methods for planning and control. Then, we shift our focus to more recent learning-based approaches, especially those leveraging reinforcement learning, imitation learning, and foundation models.
@@
Learning-based approaches have witnessed a rapid surge in humanoid robotics and achieved impressive results that attract an increasing number of researchers to the field. Among the diverse learning approaches, Reinforcement Learning (RL) has proven its ability to achieve robust motor skills. However, despite its ability to discover novel behaviors via trial and error, pure RL without demonstration data is often prohibitively inefficient for HLM tasks, which are characterized by high degrees-of-freedom robots and sparse reward settings. Therefore, training RL in simulation and transferring to the real world has become the prevalent method, though it faces the challenge of bridging sim-to-real gaps. However, imitation learning (IL) from expert demonstrations has proven to be an efficient method of acquiring motor skills. IL techniques such as behavior cloning [8] have shown impressive abilities to mimic a wide range of skills. In pursuit of versatile and generalizable policies through IL, many researchers and companies have focused on scaling data. Whereas robot experience data can be diverse and high-quality, its acquisition is both expensive and time-consuming. Thus, learning from human data, which is abundant and readily available from Internet videos and public datasets, emerges as a pivotal strategy for humanoid robotics. Learning from humans is a unique advantage exclusive to humanoid robots. However, even though humanoid robots may attain human-level motor skills, a deeper question of embodied intelligence persists: how to learn the intentions (source) behind human actions rather than merely replicating the observed motions (outcome). It is hypothesized that understanding human intention is achieved via Foundation Models (FMs) capable of semantic interpretation of the environment and the task. This hypothesis motivates us to include FMs as a part of our survey.
@@
The remarkable success of FMs has sparked a surge in general robotics research, as highlighted by several comprehensive surveys [9, 10]. This paper reports on the application of FMs for humanoid robots. FMs offer a promising solution to the persistent challenge of generalizability in robotics by efficiently harnessing internet-scale datasets to acquire extensive knowledge. A pre-trained FM exhibits capabilities for open-world reasoning and multimodal semantic understanding. These capabilities are invaluable for robots engaged in complex physical environments requiring long-term, logically coherent task planning. Within the realm of humanoid robots, FMs have been successfully implemented as task planning modules in hierarchical planning and control frameworks. However, FMs have not yet achieved robust execution of low-level sensorimotor skills in an end-to-end process. Despite the limited number of works dedicated to FMs for humanoid applications, this field is increasingly active and poised for significant future developments.
I-A Survey Goals and Roadmap
@@
This survey is related to survey papers on humanoid robotics [11, 12], and to the topics of model-based planning [13] and control [7], and learning-based methods [14]. Distinct from these survey papers, which focus on particular sub-fields of humanoid robots, our paper aims to present a broader overview, covering planning, control, and learning topics, including RL, IL, and foundation models. These provide new insights, augment the loco-manipulation capabilities enabled by traditional model-based approaches, and meet the current trends of humanoid robots within both academia and industry. In particular, we ask the following questions:
  • Q1: What are the state-of-the-art methods, both model-based and learning-based, that have already achieved loco-manipulation skills on humanoid robots?
  • Q2: What gaps still exist in achieving versatile and generalizable humanoid robots?
  • Q3: What methods are promising for addressing these gaps?
We then examine the state-of-the-art learning-based algorithms. In Sec. VII, we explore approaches using reinforcement learning and imitation learning to acquire loco-manipulation skills. In Sec. VIII, we discuss how foundation models become the backbone of semantic understanding and decision-making for effective humanoid task planning. Finally, we highlight significant challenges in this field and present our perspectives on potential future research directions and emerging opportunities in Sec. IX.
@@ II Background
In this section, we discuss the level of anthropomorphism of humanoid robots. We then focus on its main capabilities: bipedal locomotion and whole-body manipulation. Finally, we detail the combined loco-manipulation skills with the state-of-the-art methods and current challenges.
II-A Humanoid Robots
A humanoid robot refers to any anthropomorphic robot that resembles the form of a human [11]. Typically, a humanoid robot possesses a torso, two arms, and two legs, though the degree of anthropomorphism may vary. For instance, some humanoid robots feature simple hands or wheeled legs [15]. The level of anthropomorphism can be evaluated in terms of differences (with the human) in weight, limbs’ sizes, and the degrees of freedom in all joints.
The primary focus of this review is on humanoid robots that emulate human morphologies and functionalities, rather than those closely mimicking human visual appearance and external look. Because of their similar morphology to humans, humanoid robots could, in principle, exploit the abundant data that can be collected from human demonstration. In this sense, human skills are more conveniently transformable to humanoid robots. By means of scaling data and computation, humanoid robots will be more capable of versatile and generalizable skills.
From the perspective of human-robot interaction, humanoid robots would be more favorably received. This is because their human-like behavior confers in them more trust in usability as a humanoid would generate motions in an expected and predictable way by human users. This facilitates psychological comfort and also promotes effective collaboration between humans and robots, particularly in close-contact interaction tasks. Besides, humanoid robots are well-suited for environments designed for humans.
II-B Bipedal Locomotion and Navigation
Bipedal locomotion: Bipedal locomotion is a significant characteristic of humanoid robots. Therefore, in the past three decades, bipedal locomotion has been a prolific field of research in the humanoid domain. Interested readers can refer to the excellent reviews (most of which are recent) [16, 17, 18], and the monographs [19, 20]. In summary, model-based bipedal locomotion has evolved significantly, progressing from passive walking [21, 22] to quasi-static walking [23], and then to dynamic walking [19]. Bipedal walking on flat surfaces has been well explored and mastered through periodic motions with model-based methods [19, 24]. These approaches have also expanded to more agile motions such as jumping [25, 26] and back-flipping [27].
Bipedal locomotion under external perturbations and force loads has been extensively studied. Such capabilities lay the foundation for simultaneous locomotion and manipulation, the focus of this survey. Model-based approaches, such as those in [28, 29, 2, 30], have been developed to achieve such capabilities. For example, a passivity-based controller with task space dynamics is introduced in [29], where external forces are integrated as part of the generalized forces that describe the robot’s dynamics. Payload is incorporated into a simplified rigid body model to enable dynamic walking while carrying in [2]. In [28], any external force is incorporated as part of the LIP MPC.
@@ II-B Bipedal Locomotion and Navigation
Bipedal Navigation: Proficiency in bipedal locomotion has naturally progressed to advancing humanoid robots’ ability to effectively navigate complex environments, including indoor and outdoor areas with uneven terrain and dynamic obstacles. A navigation stack often incorporates a hierarchical structure: a global path planner and a local step planner. The global path planner [40, 41, 42, 43, 44] is typically responsible for understanding the overall navigation task and generating a path that avoids obstacles and reaches the target location. On the other hand, local step planners, e.g., [45, 46, 47] focus on determining the precise foot placements that adhere to the bipedal dynamics within the immediate surroundings of the robot while also tracking the global path.
From the aforementioned navigation stack, bipedal navigation capabilities have progressed from static obstacle avoidance on flat terrain [48] to more challenging scenarios, including locomotion through height-constrained space [43, 49], avoiding dynamic obstacles in a constrained environment [50], navigating dynamic social environments [51], and traversing rough terrains [40, 42, 52, 41, 53, 54]. A persistent challenge for these methods is that they are tailored to specific use case scenarios and lack the versatility to handle a wide range of different situations.
While bipedal locomotion and navigation have been widely studied, real-world deployment remains a significant challenge due to inherent uncertainties. Uncertainty can arise from the environment and the robot model. Real-world environments have uneven, varying terrain, dynamic obstacles, and occlusion, making it difficult to ensure the safety and robustness of bipedal navigation. On the other hand, model uncertainty arises from discrepancies in the mathematical representation of the robot model and the physical system. Model uncertainty also exists in most current navigation frameworks that employ reduced-ordered models at the high level for collision avoidance and goal-reaching tasks and a full-order model at the low level for tracking high-level commands. A coupled framework that considers both the navigation task and whole-body control stability and accuracy is still under-explored. Although previous works have addressed various aspects of environment uncertainties [55] and model uncertainties [56], a comprehensive navigation stack capable of handling the full spectrum of real-world uncertainties is still essential.
@@ II-D Loco-manipulation
One of the key features of humanoid robots is their ability to simultaneously perform locomotion and manipulation (abbreviated as loco-manipulation hereafter) tasks. As suggested by its name, loco-manipulation involves both the movement of objects through manipulation and the mobility of the robot self through locomotion. In a more general case of whole-body loco-manipulation, the whole-body refers to the use of all body surfaces to interact with the environment. We summarize the relationship between loco-manipulation, and whole-body manipulation in Table I. Both whole-body manipulation and loco-manipulation highlight the importance of utilizing physical contact. As shown in Fig. 3, loco-manipulation considers the movement of the robot itself while it manipulates an object, whereas whole-body manipulation emphasizes leveraging all accessible robot contact surfaces, such as using the chest as extra support to move large objects.
@@
Loco-manipulation for humanoid robots is particularly challenging, compared with quadrupeds. Humanoid robots have a smaller support region on the ground and a higher center of mass, which is challenging for dynamic balance. Therefore, early humanoid frameworks focus on separate control for locomotion and manipulation. For example, in locomotion tasks, most studies constrain the upper body to remain upright, which simplifies the whole-body problem to a bipedal locomotion problem that considers only the low limbs. Conversely, in most table-top manipulation tasks, the lower body of the humanoid remains stationary [78, 79]. In such cases, any external force exerted on the upper body is treated as a disturbance to the legs, whose goal is to solely maintain balance. On the contrary, in [1], there is no such categorization of contacts: all contacts contribute simultaneously to achieve the task and balance.
@@
Discussion: From the planning and control perspective, should we design a unified framework for humanoid robots to achieve agile locomotion and dexterous manipulation simultaneously, or treat them as separate problems in a hierarchical framework? A unified framework generates coherent whole-body motions, similar to how humans move and manipulate objects. The unified framework would allow simultaneous optimization of locomotion and manipulation, adapting to a wider range of tasks without needing to switch between modes. Considering a hierarchical framework might seem a modular solution since each layer can be optimized independently, the overall framework accommodates new tasks or modifications easily. The main downside is the lack of mutual awareness between layers. For example, if the locomotion layer does not account for manipulation needs (e.g., positioning the robot for optimal reach), the overall performance may be suboptimal.

### v2

@@ Abstract
Humanoid robots hold great potential to perform various human-level skills, involving unified locomotion and manipulation in real-world settings. Driven by advances in machine learning and the strength of existing model-based approaches, these capabilities have progressed rapidly, but often separately. This survey offers a comprehensive overview of the state-of-the-art in humanoid locomotion and manipulation (HLM), with a focus on control, planning, and learning methods. We first review the model-based methods that have been the backbone of humanoid robotics for the past three decades. We discuss contact planning, motion planning, and whole-body control, highlighting the trade-offs between model fidelity and computational efficiency. Then the focus is shifted to examine emerging learning-based methods, with an emphasis on reinforcement and imitation learning that enhance the robustness and versatility of loco-manipulation skills. Furthermore, we assess the potential of integrating foundation models with humanoid embodiments to enable the development of generalist humanoid agents. This survey also highlights the emerging role of tactile sensing, particularly whole-body tactile feedback, as a crucial modality for handling contact-rich interactions. Finally, we compare the strengths and limitations of model-based and learning-based paradigms from multiple perspectives, such as robustness, computational efficiency, versatility, and generalizability, and suggest potential solutions to existing challenges.
@@ I Introduction
Figure 1: Humanoids executing locomotion and manipulation tasks: (a) HRP-4 wipes a wood board while adapting to terrain [1]; (b-g) Object pick and place by Digit, Hector [2], Atlas, H1, TORO [3], and Apollo; (h) iCub pushes a cart [4]; (i) Nadia opens a door [5]; (j-k) Object manipulation by Figure 02 and Optimus; (l) MIT humanoid whole-body push recovery [6].
A humanoid robot refers to any anthropomorphic robot that resembles the form of a human [7]. Typically, a humanoid robot possesses a torso, two arms, and two legs, though the degree of anthropomorphism may vary. For instance, some humanoid robots feature simple hands or wheeled legs [8]. The primary focus of this review is on humanoid robots that replicate human morphologies and functionalities, rather than those that closely mimic human-like visual appearance and external aesthetics. Specifically, this survey examines bipedal humanoid robots, with an emphasis on the whole-body motion of humanoid robots instead of on the dexterous manipulation of multi-finger hands.
Humanoid robots are well suited for executing human-level tasks, as they are built to (ideally) replicate human motions and achieve various whole-body loco-manipulation tasks, e.g., applications ranging from manufacturing to services, as shown in Fig. 1. Their anthropomorphism makes them stand out from other robot forms in performing human-like tasks, such as payload carrying over stairs, especially in environments designed for humans. Furthermore, humanoid robots can interact with humans for physical collaboration tasks, such as collaboratively moving a large table upstairs with human assistance. However, simultaneously achieving these intricate tasks while addressing highly complex robot dynamics is still challenging, let alone safe physical collaborations with humans and/or operations in unstructured environments. As a promising direction to solve this problem, humanoids can exploit the abundance of data collected from humans to quickly acquire motor and cognitive skills due to their human-like morphology. Therefore, scaling human data and computational resources for humanoid embodiment is potentially a fast route to embodied intelligence.
Embodied intelligence at large is thriving at an unprecedented pace. Perception algorithms can detect, classify, and segment a wide variety of objects in real time. Model-based methods that leverage both predictive and reactive control have enabled agile and reliable locomotion and manipulation. Meanwhile, deep learning policies have demonstrated convincing control results on robot hardware through exploration and imitation. Large foundation models trained on massive, internet-scale datasets are beginning to show cognitive capabilities of open-world reasoning. Consequently, the building of autonomous humanoid robots for real-world applications has become possible, leading to the emergence of many humanoid robot companies and concrete deployment applications. Especially, companies with powerful GPU-based parallelization capabilities, such as NVIDIA, and companies building physical humanoid robots, such as Boston Dynamics, Tesla, and Figure, have begun collaboration on the embodied intelligence of humanoid robots.
@@
Acknowledging the rapid advancements in humanoid robotics, this article reviews recent developments in Humanoid Locomotion and Manipulation (HLM). As laid out in Fig. 2, humanoid robotics is a multidisciplinary field that spans domains in design, actuation, sensing, control, planning, and decision-making. In this survey, we mainly examine task planning, motion planning, policy learning, and control from the perspective of model-based and learning-based methods. We first synthesize traditional model-based methods for planning and control. Then, we shift our focus to more recent learning-based approaches, especially those leveraging reinforcement learning, imitation learning, and foundation models.
@@
Learning-based approaches have witnessed a rapid surge in humanoid robotics and achieved impressive results that attract an increasing number of researchers to the field. Among the diverse learning approaches, Reinforcement Learning (RL) has proven its ability to achieve robust motor skills. However, despite its ability to discover novel behaviors via trial and error, pure RL without demonstration data is often prohibitively inefficient for HLM tasks, which are characterized by high degrees-of-freedom robots and sparse reward settings. Therefore, sim-to-real RL, where training is conducted in simulation before transferring to the physical robot, has become a prevalent method, though it faces the challenge of bridging sim-to-real gaps. From the existing literature, we observe that sim-to-real RL has gained significant traction for achieving humanoid whole-body motion. There is an ongoing debate between model-based and learning-based control, and therefore we highlight two key points: (1) sim-to-real RL that relies on the accurate dynamics models in a simulator can also be viewed as a “model-based,” method, even though the RL algorithm itself does not explicitly model the dynamics; and (2) model-based control and sim-to-real RL do not conflict; on the contrary, they often complement each other and can be combined to achieve a better performance than either method alone.
@@
Imitation learning (IL) from expert demonstrations has proven to be an efficient method of acquiring motor skills. IL techniques, such as behavior cloning [10], have shown impressive abilities to mimic a wide range of skills. In pursuit of versatile and generalizable policies through IL, many researchers and companies have focused on scaling data. While robot experience data can be diverse and high-quality, its acquisition is both expensive and time-consuming. Thus, learning from human data, which is abundant and readily available from Internet videos and public datasets, emerges as a pivotal strategy for humanoid robotics. Learning from humans is a unique advantage exclusive to humanoid robots. However, even though humanoid robots may attain human-level motor skills, a deeper question of embodied intelligence persists: how to learn the intentions (source) behind human actions rather than merely replicating the observed motions (outcome). It is hypothesized that understanding human intention is achieved via Foundation Models (FMs) that are capable of semantic interpretation of the environment and the task. This hypothesis motivates us to include FMs as a part of our survey.
@@
The remarkable success of FMs has sparked a surge in general robotics research, as highlighted by several comprehensive surveys [11, 12]. This paper surveys the application of FMs specifically for humanoid robots. FMs offer a promising solution to the persistent challenge of generalizability in robotics by efficiently harnessing internet-scale datasets to acquire extensive knowledge. A pre-trained FM exhibits capabilities for open-world reasoning and multimodal semantic understanding. These capabilities are invaluable for robots engaged in complex environments requiring long-term, logically coherent task planning. Within the realm of humanoid robots, FMs have been successfully implemented as task planning modules in hierarchical planning and control frameworks. Training end-to-end FMs for humanoid applications is an emerging area, poised for significant future advancements.
I-A Survey Roadmap
For the selection criteria of papers cited in this survey, we prioritize topics most relevant to humanoid loco-manipulation capabilities, including tactile sensing [13], model-based planning [14] and control [9], as well as learning-based methods such as RL [15, 16], IL [17, 18], and FMs [12, 11]. Distinct from survey papers on these topics, our paper aims to present a broad overview with a focus on humanoid loco-manipulation.
Each of these topics has extensive studies. For each topic, we aim to spotlight three categories of papers: (i) foundational studies that established key principles in these areas, (ii) recent works representing the state of the art, and (iii) survey papers for further reading. Our survey spans 30 years of research. Specifically, model-based planning and control papers have been collected since 1990, and learning-based methods predominantly span the past 20 years.
Compared with existing survey reviews on humanoid robotics [7, 19], this survey collects up-to-date papers and introduces in detail the new paradigm of learning-based methods.
We then examine the state-of-the-art learning-based algorithms. In Sec. VII, we explore approaches using reinforcement learning and imitation learning to acquire loco-manipulation skills. In Sec. VIII, we discuss how foundation models have become the backbone of semantic understanding and decision-making for effective humanoid task planning. Finally, we highlight significant challenges in this field and present our perspectives on potential future research directions and emerging opportunities in Sec. IX.
I-B Survey Contributions
We summarize our core contributions as follows:
@@ I-B Survey Contributions
  • We advance the understanding of loco-manipulation in bipedal humanoid robots by clearly defining the problem, surveying state-of-the-art methods and capabilities, and outlining remaining challenges along with promising solutions.
  • We provide a clear comparison between model-based and learning-based methods, and their corresponding loco-manipulation skills achieved on humanoid robots. This comparison offers valuable insights for researchers deciding their methods.
@@ II Background
In this section, we focus on the humanoid robot’s main capabilities: bipedal locomotion and whole-body manipulation. Finally, we detail the combined loco-manipulation skills with the state-of-the-art methods and current challenges.
II-A Bipedal Locomotion
Bipedal locomotion is a significant characteristic of humanoid robots. Therefore, in the past three decades, bipedal locomotion has been a prolific field of humanoid research. Interested readers can refer to the excellent reviews (most of which are recent) [20, 21, 22], and the monographs [23, 24]. In summary, model-based bipedal locomotion has evolved significantly, progressing from passive walking [25, 26] to quasi-static walking [27], and then to dynamic walking [23]. Bipedal walking on flat surfaces has been well explored and mastered through periodic motions with model-based methods [23, 28]. These approaches have also expanded to more agile motions such as jumping [29, 30] and back-flipping [31].
Bipedal locomotion under external perturbations and force loads has been extensively studied. Such capabilities lay the foundation for simultaneous locomotion and manipulation, the focus of this survey. Model-based methods are among the early works that demonstrate these capabilities. For example,  [32] introduced a passivity-based controller with task space dynamics that treats external forces as part of the robot’s generalized forces. [2] incorporated a payload into a simplified rigid body model to enable dynamic walking while carrying a load, and [33] integrated external forces into a simplified linear inverted pendulum model.
@@ II-A Bipedal Locomotion
Researchers have explored energy efficiency in bipedal locomotion by examining the Cost of Transport (COT), which designates energy expenditure per unit distance traveled, normalized by body weight. Early work in passive dynamic walking [25, 26] achieved higher efficiency than humans but offered limited versatility. Meanwhile, today’s humanoid robots $(\text{COT}>0.7)$ are significantly less efficient than humans $(\text{COT}=0.2)$ [43]. Therefore, efficient walking continues to motivate hardware designs that incorporate passive-compliant components for energy storage [44, 45] and control algorithms that exploit the system’s natural dynamics [46, 47].
II-B Bipedal Navigation
Proficiency in bipedal locomotion has naturally progressed to humanoid robots’ ability to effectively navigate complex environments, including indoor and outdoor areas with uneven terrain and dynamic obstacles. A navigation stack often incorporates a hierarchical structure: a global path planner and a local step planner. The global path planner [48, 49, 50, 51, 52] is typically responsible for understanding the overall navigation task and generating a path that avoids obstacles and reaches the target location. On the other hand, local step planners, e.g., [53, 54, 55] focus on determining the precise foot placements that adhere to the bipedal dynamics within the immediate surroundings of the robot while also tracking the global path.
From the aforementioned navigation stack, bipedal navigation capabilities have progressed from static obstacle avoidance on flat terrain [56] to more challenging scenarios, including locomotion through height-constrained space [51, 57], avoiding dynamic obstacles in a constrained environment [58], navigating dynamic social environments using reachable sets [59], or deep RL that incorporates the perceived pedestrian emotions into the navigation plan [60], and traversing rough terrains [48, 50, 61, 49, 62, 63]. A persistent challenge for these methods is that they are tailored to specific use case scenarios and lack the versatility to handle a wide range of different situations.
While bipedal locomotion and navigation have been widely studied, real-world deployment remains a significant challenge due to inherent uncertainties. Uncertainty can arise from the environment and the robot model. Real-world environments have uneven, varying terrain, dynamic obstacles, and occlusion, making it difficult to ensure the safety and robustness of bipedal navigation. On the other hand, model uncertainty arises from discrepancies in the mathematical representation of the robot model and the physical system. Model uncertainty also exists in most current navigation frameworks that employ reduced-ordered models at the high level for collision avoidance and goal-reaching tasks and a full-order model at the low level for tracking high-level commands. A coupled framework that considers both the navigation task and whole-body control stability and accuracy is under-explored. Although previous works have addressed various aspects of environment uncertainties [64] and model uncertainties [65], a comprehensive navigation stack capable of handling the full spectrum of real-world uncertainties is still essential.
@@ II-D Loco-manipulation
One of the key features of humanoid robots is their ability to simultaneously perform locomotion and manipulation (abbreviated as loco-manipulation hereafter) tasks. As suggested by its name, loco-manipulation involves both the movement of objects through manipulation and the mobility of the robot itself through locomotion. In a more general case of whole-body loco-manipulation, the whole-body refers to the use of all body surfaces to interact with the environment. We summarize the relationship between loco-manipulation, and whole-body manipulation in Table I. Both whole-body manipulation and loco-manipulation highlight the importance of utilizing physical contact. As shown in Fig. 3, loco-manipulation considers the movement of the robot itself while it manipulates an object, whereas whole-body manipulation emphasizes leveraging all accessible robot contact surfaces, such as using the chest as extra support to move large objects.
@@
Loco-manipulation for humanoid robots is particularly challenging, compared with quadrupeds. Humanoid robots have a smaller support region on the ground and a higher center of mass, which is challenging for dynamic balance. Therefore, early humanoid frameworks focus on separate control for locomotion and manipulation. For example, in locomotion tasks, most studies constrain the upper body to remain upright, which simplifies the whole-body problem to a bipedal locomotion problem that considers only the lower limbs. Conversely, in most table-top manipulation tasks, the lower body of the humanoid remains stationary [87, 88]. In such cases, any external force exerted on the upper body is treated as a disturbance to the legs, whose goal is to solely maintain balance. On the contrary, in [1], there is no such categorization of contacts: all contacts contribute simultaneously to achieve the task and balance.